What would be the pressure of a gas, at 101.3 kPa, if the temperature is lowered from 368 K to 358 K?
1
Expert's answer
2020-05-19T08:46:45-0400

According to Gay-Lussac's Law:

p1/T1 = p2/T2

where p1 - initial pressure, T1 - initial temperature, p2 - final pressure, T2 - final temperature.

From here:

p2 = p1 × T2/T1 = 101.3 kPa × 358 K / 368 K = 98.5 kPa


Answer: 98.5 kPa